OpenStudy (valiyuh):

-14+ 3 ( x + 10 ) = 7 ( 2x + 4 ) + x I need help solving this algebra problem.

OpenStudy (lynfran):

here an example same structure -2+5(x+3)=3(3x+1)+x -2+5x+15=9x+3+x 5x+13=10x+3 5x+13-3=10x 13-3=10x-5x 10=5x 10/5=x 2=x

OpenStudy (anonymous):

Did you understand her example? That pretty much it, but the coefficients are different.

OpenStudy (valiyuh):

Oh alright i see it now thank you so much !☺ @LynFran Yes i did understand it thank you! @mathway